OpenStudy (anonymous):

Look at the quadratic equation below. -x2 + 3 = (2x - 1)2 What is this quadratic equation written in standard form? A. 5x2 - 4x - 2 = 0 B. 3x2 - 4x - 2 = 0 C. 3x2 - 4x + 4 = 0 D. 5x2 - 4x + 4 = 0

OpenStudy (anonymous):

@pooja195 @jigglypuff314 @Michele_Laino @imqwerty @KendrickLamar2014 @rebeccaxhawaii

imqwerty (imqwerty):

\(-x^2 + 3 = (2x - 1)^2 \) step 1- simplify the right hand side part to do that use this identity- \((a-b)^2=a^2+b^2-2ab\) step2- bring all the terms one one side step 3- combine like terms and simplify

OpenStudy (anonymous):

can u do it with me

OpenStudy (anonymous):

how do i simplfiy the right side

imqwerty (imqwerty):

yeah :0

imqwerty (imqwerty):

* :)

OpenStudy (anonymous):

how do i do \((a-b)^2=a^2+b^2-2ab\)

imqwerty (imqwerty):

okay we go like this- we use this identity- \((a-b)^2=a^2+b^2-2ab\) now we look at our equation- \(-x^2 + 3 =\color{red}{ (2x - 1)^2}\) simplifying- \(-x^2+3=(2x)^2 +(1)^2 -2(2x)(1)\)

OpenStudy (anonymous):

confused ): that stuff looks really complicated

imqwerty (imqwerty):

its okay :) did you get that part where we used the identity?

OpenStudy (anonymous):

what is with the colors?

OpenStudy (anonymous):

maybe if u drew it i would understand? because u dont get that );

imqwerty (imqwerty):

i just used the colors to co relate that bracket thing with the identity

OpenStudy (anonymous):

is there another way of doing this? or another way to show me

imqwerty (imqwerty):

okay what we did was- \(-x^2 + 3 =\color{red}{ (\color{blue}{2x} - \color{green}1)^2}\) we need to expand that bracket on the right so we use this identity-> \((a-b)^2=a^2+b^2-2ab\) look at the colored thing \(\color{red}{(\color{blue}{2x} - \color{green}1)^2}\) now use this identity-> \(\color{red}{(\color{blue}a-\color{green}b)^2}=a^2+b^2-2ab\) \(\color{red}{(\color{blue}{2x} - \color{green}1)^2}=(2x)^2+(1)^2-2(2x)(1)\) so our equation becomes- \(-x^2 + 3 =\color{red}{ (\color{blue}{2x} - \color{green}1)^2}\) \(-x^2 + 3 =(2x)^2+(1)^2-2(2x)(1)\)

imqwerty (imqwerty):

maybe this could help- http://prntscr.com/amjf6m

OpenStudy (anonymous):

prntscr is blocked on school internet

OpenStudy (anonymous):

i looked it up on my phone can u explain the picture

imqwerty (imqwerty):



OpenStudy (anonymous):

3-x²=(2x-1)²

imqwerty (imqwerty):

are you stuck on the 2nd step?

OpenStudy (anonymous):

yes

imqwerty (imqwerty):

are you familiar with this identity-> \((a-b)^2=a^2+b^2-2ab\)

OpenStudy (anonymous):

no, i am fimilar with ax²+bx+c=0

imqwerty (imqwerty):

well in this question we need to use that identity so i suggest that you watch this video-> https://youtu.be/Gr1E7TtXQ9U it will help you clear that 2nd step

OpenStudy (anonymous):

youtube is blocked on school internet

imqwerty (imqwerty):

well then check out the 2nd identity given here- http://www.ask-math.com/factorization-using-identities.html

OpenStudy (anonymous):

1) a2 + 2ab + b2 = (a + b)2 2) a2 – 2ab + b2 = (a – b)2 3) a2 – b2 = (a + b) (a – b) 4) x2 + (a + b) x + ab = (x + a) (x + b)